OpenStudy (anonymous):

Mandy was laying tile in her kitchen. She needed 2 and five-eighths rows with 3 and one-half tiles per row to cover the entire floor. How many total tiles did she need? Multiply to find out. 6 and five-sixteenths 6 6 and one-half 9 and three-sixteenths

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@student_basil

OpenStudy (anonymous):

6 1/2

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

\(\Large\color{blue}{ \bf 2 \frac{5 }{8} \times 3 \frac{1 }{2}}\) \(\Large\color{blue}{ \bf \frac{21 }{8} \times \frac{7 }{2}}\) \(\Large\color{blue}{ \bf \frac{147 }{16}}\) \(\Large\color{blue}{ \bf \frac{144+3}{16}}\) \(\Large\color{blue}{ \bf 9 \frac{3}{16}}\)

OpenStudy (anonymous):

:)